Operations Report – December 22nd

Crew 236 Operations Report 22Dec2021

SOL: 3

Name of person filing report: Pavithra Ravi

Non-nominal systems: Suits and Radios — a headpiece is malfunctioning

Notes on non-nominal systems: One of the headsets is malfunctioning. It keeps transmitting even when the earpiece button is not pressed, resulting in loud interference. It was verified that the issue is with the headset and not the radio (the issue persisted when the radio was swapped out).

Generator:

Hours run: 13 hours 18 minutes

From what time last night: 18:27

To what time this morning: 07:45

List any additional daytime hours when the generator was run: None

Solar— SOC % (Before generator is run at night): Information provided by Outpost

Station Propane Reading – % Information provided by Outpost

Ethanol Free Gasoline: Not in use

Water (loft tank): 54 gallons

Water Meter: 1539651 units

Water (static tank): approx. 370 gallons

Static to Loft Pump used: Yes. The loft tank was refilled from 12.5 to 54 gallons.

Water in GreenHab: approx. 222 gallons

Water in ScienceDome: N/A

Toilet tank emptied: No

Deimos rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No

Hours (before EVA 1): N/A

Beginning charge (Before EVA 1): N/A

Ending charge (On return from EVA 1, before recharging): N/A

Hours (before EVA 2): N/A

Beginning charge (Before EVA 2): N/A

Ending charge (On return from EVA 2, before recharging): N/A

Currently charging: Yes

Opportunity rover used: Yes

Hours: 76.4

Beginning charge: 100%

Ending charge: 99%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 154.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: None

ATV’s Used (Honda, 300, 350.1, 350.2, 350.3): None

Reason for use: N/A

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were used today: N/A

Notes on ATVs: N/A

HabCar used and why, where? No

CrewCar used and why, where? No

General notes and comments: None

Summary of internet: Nominal

Summary of suits and radios: One of the headsets is malfunctioning. It keeps transmitting even when the earpiece button is not pressed, resulting in loud interference. It was verified that the issue is with the headset and not the radio (the issue persisted when the radio was swapped out).

Summary of Hab operations: Nominal

Summary of GreenHab operations: Nominal

Summary of ScienceDome operations: Nominal

Summary of RAM operations: Nominal

Summary of any observatory issues: In correspondence with the
astronomy support team.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support:

The heater filter was changed today. Should the old filter be added to the non-burnables, or be placed in the back airlock for removal?

Any suggestions on how to fix the earpiece? If it is not fixable,
could a new earpiece be provided so each crew member can have their own?